Need help solving assignment due tongiht.
2) Use the geometric sequence of numbers 1, 3, 9, 27, … to find the following:
a) What is r, the ratio between 2 consecutive terms?
Answer:
Show work in this space.



b) Using the formula for the nth term of a geometric sequence, what is the 10th term?
Answer:
Show work in this space.



c) Using the formula for the sum of a geometric series, what is the sum of the first 10 terms?
Answer:
Show work in this space.

Answer by venugopalramana(3286)

SEE THE FOLLOWING SIMILAR EXAMPLES AND TRY.IF STILL IN DIFFICULTY PLEASE COME BACK
-------------------------------------------------------
1) Use the geometric series
of numbers 1, 2, 4, 8,…to find the following:
a) What is r, the ratio between 2 consecutive terms?
b) Using the formula for the nth term of a geometric
series, what is the 24th term?
c) Using the formula for the sum of a geometric
series, what is the sum of the first 10 terms?
pls help
1 solutions
Answer 14006 by venugopalramana(1619)on
2006-01-31 05:56:57 
) Use the geometric series of numbers 1, 2, 4, 8,…to
find the following:
a) What is r, the ratio between 2 consecutive terms?
b) Using the formula for the nth term of a geometric
series, what is the 24th term?
c) Using the formula for the sum of a geometric
series, what is the sum of the first 10 terms? SUM IS
GIVEN BY
SN=A(R^N-1)/(R-1)
S10=1*(2^10-1)/(2-1)=2^10-1
SEE THE FOLLOWING EXAMPLE AND SOLVE
Use the geometric series of numbers 1, 1/2, 1/4,
1/8,…to find the following:
a) What is r, the ratio between 2 consecutive terms?
R=(1/2)/1=1/2...IT IS SAME FOR ANY 2 CONSECUTIVE
TERMS...SAY
(1/8)/(1/4)=1/2...ETC......
b) Using the formula for the nth term of a geometric
series, what is 10th term?
TN=A*(R)^(N-1)=1*(1/2)^(N-1)
T10=(1/2)^(10-1)=(1/2)^9
c) Using the formula for the nth term of a geometric
series, what is 12th term?
T12=(1/2)^11
d) What observation can make about these sums? In
particular, what number does it appear that the sum
will always be smaller than?
THE NUMBERS TEND TO ZERO AS N INCREASES TO LARGE
NUMBERS AND TOWARDS INFINITY.
SUM OF A G.P. TO INFINITE TERMS WITH R<1 IS GIVEN BY
A/(1-R)=1/(1-0.5)=2
Geometric_formulas/26075: 10) Use the geometric series
of numbers 1, 1/2, 1/4, 1/8,…to find the following:
a) What is r, the ratio between 2 consecutive terms?
b) Using the formula for the nth term of a geometric
series, what is 10th term?
c) Using the formula for the nth term of a geometric
series, what is 12th term?
d) What observation can make about these sums? In
particular, what number does it appear that the sum
will always be smaller than?
1 solutions
Answer 14005 by venugopalramana(1619)on
2006-01-31 05:54:54 
Use the geometric series of numbers 1, 1/2, 1/4,
1/8,…to find the following:
a) What is r, the ratio between 2 consecutive terms?
R=(1/2)/1=1/2...IT IS SAME FOR ANY 2 CONSECUTIVE
TERMS...SAY
(1/8)/(1/4)=1/2...ETC......
b) Using the formula for the nth term of a geometric
series, what is 10th term?
TN=A*(R)^(N-1)=1*(1/2)^(N-1)
T10=(1/2)^(10-1)=(1/2)^9
c) Using the formula for the nth term of a geometric
series, what is 12th term?
T12=(1/2)^11
d) What observation can make about these sums? In
particular, what number does it appear that the sum
will always be smaller than?
THE NUMBERS TEND TO ZERO AS N INCREASES TO LARGE
NUMBERS AND TOWARDS INFINITY.
SUM OF A G.P. TO INFINITE TERMS WITH R<1 IS GIVEN BY
A/(1-R)=1/(1-0.5)=2